Blog Archive

Last modified by Michael Sibayan on 2015/02/13 12:22
Comments (0) · Attachments (0) · History

Blog posts for November 2016

Irrigation valve actuator swap 2016117093157

On Nov. 6, 2016 from about 10:45 PM - 11:30 PM Aaron and Michael performed work on the electric actuators controlling/monitoring LEO irrigation. 

Several of the actuators were not responding (No diagnostics have been performed), so we swapped around actuators such that each bay now should have a functional actuator on both the MULTI1 and MULTI2 circuit. 

Due to this change several irrigation circuits will display as "in transition" because they have been physically removed / electrically disconnected.

LEO bigPERTH experiment started

The LEO bigPERTH experiment was started on Nov. 6, 2016 at 23:00 MST.

Irrigation actuators on LEO East and West did not respond, so circuits were opened and closed in manual mode with a slight start-time delay (1-5 minutes).

20161110100054 XWiki.eahunt

Vacuum started at 0930 on all samplers for all three hill slopes.  Vacuum stopped at 0630 next morning.  Collected all sample available as Set 1 for geochemistry.

bigPERTH_SoilCoring_Beginning of Experiment

Soil cores were collected from all three hillslopes on 11/8/2016.

  • See attached document for core location and nomenclature
  • 6 cores per slope * 3 slopes = 18 cores

    **Once coring is done, fill gaps with blue barrel soil material that have received the same volume of total water as the big hillslope.SoilCore_bigPERTH.jpgSoilCore_bigPERTH.jpg

    Coring (Nate, Aaron); LEO spaceProcessing (Ed/Aditi); Analytical LabStoring/Archiving (Ed/Aditi)

    1.    Core at 6 locations per slope

    2.    Tape core ends shut with electrical tape, 

    label cores with labeling tape to indicate surface and deepest end. 

    Store cores in cooler prior to processing

    3.    Hand over cores to Ed/Aditi

    4.     Measure cores for sub-sections

    5.     Cut cores in the lab

    6.     Place in labeled bags, homogenize

    7.     Aliquot for DNA, RNA in microcentrifuge tubes (plus archive in 50 ml tubes),

    8. Flash freeze microbio samples, drop in dry ice till stored at -80

    9. Ed takes care of geochem samples

  • Update on sampling:
  • Central slope was cored first,starting at approximately 10.45 AM-12.00PM. Followed by west slope (starting at 1.30 PM to 3 PM), followed by east slope (3.45-5.30 pm). Central slope cores were collected one at a time (Cores were stored on dry ice while being collected). West and East slope cores were collected as one set (Nate and Aaron did not hand down each core prior to collecting the next one. So east and west cores were not stored in cooler while being collected). The metal core has to be cleaned every time with water so for east and west cores, wash bottle was carried on the gantry along with a bucket to wash the cores. Also, location 2-6 are cored at one go, for location 1, power drill connection has to be pulled from the top of the slope. 
  • For next coring, buy big cooler, add ice, drop cores in it. Ask Nate and Aaron to carry cooler with them on gantry. 
  • The average core length was 65-70 cm. Four 15 cm sub-depths were collected from each core plus a fifth depth sample that comprised of whatever was left over.
  • Sub-samples were homogenized in whirl pack bags, Ed took C and N samples from the bags, bag handed to Aditi,  who subsampled into centrifuge tubes. Started out with sampling into 1.5 ml centrifuge tubes. Stopped after central slope. Way too much time was being spent in doing this (the integrity of the cores was in question since they were out in bags now). Rest of the samples are in whirlpack bags and 50 ml centrifuge tubes. All bags and tubes were flash frozen and stored on dry ice, then brought back to campus and stored at -80 C at Raina Maier's lab.  An alternate is to have more people work to make sure this gets done faster. Plus have another cooler of ice to store homogenized soil bags in before sub-sampling.

Seepage clogged

Seepage clogged and fixed,flow rates down while fixing ...

PE102-075 flow meter added to Center Bay Leak collector

Added a Seametrics PE102-075 flow meter inline with the Leak tipping bucket in the Center Bay. 

Valid flow rates for this model are 0.2 - 20 GPM. 

Data for Leak tipping bucket between 11/15/2016 11:40 - 13:06 are not valid

Instrumentation panel damage - West P03

At approximately 9:30 on Nov. 21 an instrumentation panel (West-P03) was damaged along with a few sensor cables. Repairs underway.

Tags:
Created by Michael Sibayan on 2015/02/13 12:22

This wiki is licensed under a Creative Commons 2.0 license
XWiki Enterprise 6.0.1 - Documentation